OpenStudy (anonymous):

What is the approximate volume of the cylinder? height = 6.4 m; diameter = 4.6 m A. 34 m3 B. 46 m3 C. 92 m3 D. 106 m3

OpenStudy (anonymous):

 \[V = hr^2\pi\]\[V = \pi(2.3)^2(6.4)\]\[V = 33.856\pi\] V ~= 106m3 D is the correct answer